OpenStudy (ellaella02):

Which answer shows this equation in standard form? 6 – 2(x – y) = –3x + 5 A. x + 2y = –1 B. x + 2y = 1 C. x – 2y = 1 D. –5x + 2y = –1

OpenStudy (ellaella02):

I'll medal!! Please help!

OpenStudy (ellaella02):

I think it's A or D but I'm not sure.

Directrix (directrix):

Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers.

OpenStudy (ellaella02):

So would it be D ? @Directrix

Directrix (directrix):

I do not know. We have to simplify the equation first.

OpenStudy (ellaella02):

6 – 2(x – y) = –3x + 5 would be 3-1(x-y)=-1.5x + 2.5x

OpenStudy (ellaella02):

Would that be right?

Directrix (directrix):

6 – 2(x – y) = –3x + 5